show that for two capacitor of area A seperated by a distance D apart with a di electric material of tickness T in between them. a capacitance is given by the expression C= permitivity of A (1/d-1 +permitivity of r/t


1
Expert's answer
2021-04-16T07:21:46-0400

"c=\dfrac{Q}{A}" (1)


"\sigma=" surface charge density

"\sigma=\dfrac{Q}{A}"

"Q=\sigma A"


"c=\dfrac{\sigma A}{V}" from (1)


Distance of plates from each other = "D"

Thickness of slab = "T"

"t_1+t_2+T=D"

"t_1+t_2=D-T"


Potential difference

"V=\dfrac{\sigma D}{\epsilon_o}"

"\Rightarrow V=\dfrac{\sigma}{\epsilon_o}(t_1+t_2+\dfrac{T}{K})"

"\Rightarrow V=\dfrac{\sigma}{\epsilon_o}(D-T+\dfrac{T}{K})"

"\Rightarrow \dfrac{Q}{c}=\dfrac{\sigma}{\epsilon_o}(D-T+\dfrac{T}{K})"

"\Rightarrow \dfrac{\sigma A}{c}=\dfrac{\sigma}{\epsilon_o}(D-T+\dfrac{T}{K})"

"\Rightarrow c=\dfrac{\epsilon_o A}{(D-T)+\dfrac{T}{K}}"